# Glauber dynamics for the hard-core model on bounded-degree $H$-free graphs

### v1

@@ 5. $S_{2,2,2}$-free graphs and beyond
First assume that $T$ has a single vertex $w$ such that either $w\not=v$ and $\deg(w)=4$ or $w=v$ and $\deg(w)=3$. Note that all vertices of $T$ other than $v,w,a,b,c$ have degree 2. By adding $e$ to $E(T)$ and removing the edges of $T$ lying between $w$ and $u_{j+1}$ we obtain a smaller tree, contradicting minimality of $T$. (Notice that we must remove at least two edges.) By choice of $a,b,c$ we know that $d_{T}(w,a),d_{T}(w,b),d_{T}(w,c)\geq t+1$, so $G$ contains an induced $S_{t,t,t}$. (Note that we are measuring graph distances within the tree $T$ here.)
Next assume that $T$ has two vertices $w$ and $w^{\prime}$, with $w^{\prime}$ in a higher numbered level than $w$, and such that either (i) $w\not=v$ and and $\deg(w)=\deg(w^{\prime})=3$ or (ii) $w=v$ and $\deg(w)=2$ and $\deg(w^{\prime})=3$. Note that all vertices of $T$ other that $v,w,w^{\prime},a,b,c$ have degree 2. Suppose, without loss of generality, that $a$ and $b$ are descendants of $w^{\prime}$. With two exceptions, we may add $e$ to $E(T)$ and remove all the edges between $u_{j+1}$ and $w$, or $w^{\prime}$ as appropriate, to obtain a smaller tree, contradicting minimality of $T$. Those exceptions occur when $w^{\prime}\in L_{j}$, i.e., $u_{j}$ is in the same level as $w^{\prime}$. In that case, one edge is added and one removed and the size of the tree remains unchanged. There are two candidate vertices for $u_{j+1}$; these are the vertices in $L_{j+1}\cap V(T)$, say $x$ and $y$, that are adjacent to $w^{\prime}$. Let $x$ (respectively, $y$) be the vertex on the path in $T$ from $w^{\prime}$ to $a$ (respectively $b$). To summarise: The graph induced by $V(T)$ is tree $T$, with the possible addition of one or two of the edges $\{u_{j},x\}$ and $\{u_{j},y\}$. (Refer to Figure 5.)
@@
By choice of $a,b,c$ we know that $d_{T}(w^{\prime},a),d_{T}(w^{\prime},b)\geq t+1$. In the case that neither of the exceptional edges in present, we have an induced $S_{t,t,t}$ with its centre at $w^{\prime}$ and with the path from $w^{\prime}$ to $c$ routing via $w$. In the case of one exceptional edge, say $\{u_{j},x\}$, we have an induced $S_{t,t,t}$ with its centre at $x$ and with the path from $x$ to $b$ routing via $w^{\prime}$, and the path from $x$ to $c$ routing via $u_{j}$. In the case of two exceptional edges, we have an induced $S_{t,t,t}$ with its centre at $u_{j}$ and with the path from $u_{j}$ to $a$ routing via $x$, and the path from $u_{j}$ to $b$ routing via $y$.
@@
For any $\Delta\geq 3$, $\lambda>0$ and $t\geq 2$, Glauber dynamics at fugacity $\lambda$ has optimal (i.e., $O(n\log n)$) mixing time on the class of $S_{t,t,t}$-free graphs of maximum degree $\Delta$.
@@ 6. Exponential-time mixing of $H$-free graphs
Let $H$ be a connected graph that is not a subdivided claw or a path. There exists a fugacity $\lambda$ and an infinite family $\mathcal{G}$ of $H$-free graphs of maximum degree 3 such that Glauber dynamics for the hard-core distribution on these graphs has exponential mixing time on $\mathcal{G}$.
@@
Let $\alpha>0$ be sufficiently small. Given $n$, let $G$ be a cubic bipartite graph on $n+n$ vertices that is a $1+\alpha$ expander for sets up to size $\frac{2}{3}n$. Such graphs exists by a lemma due to Bassalygo [5]; see Alon [3, Lemma 4.1]. Let $V_{L}\cup V_{R}=V(G)$ be the bipartition of the vertex set of $G$. Let $\ell$ be a number and let $G^{*}$ denote the $(2\ell+1)$-stretch of $G$; that is, $G^{*}$ is obtained from $G$ by subdividing each edge by $2\ell$ new vertices. Fix $\ell$ sufficiently large that $G^{*}$ contains no induced $H$. Let $V_{L}$ and $V_{R}$ continue to denote the vertices in $G^{*}$ that are inherited from $G$.
@@
We will show that the independent sets in $\mathcal{I}_{=}$ all have cardinality significantly less than $(3\ell+1)n$. So if $\lambda$ is sufficiently large, the total weight $\mathrm{wt}(\mathcal{I}_{=})$ of independent sets in $\mathcal{I}_{=}$ will be exponentially smaller than $\mathrm{wt}(\mathcal{I}_{<})$, even though the former are more numerous. (‘Energy’ wins out over ‘entropy’.) So consider an independent set $I$ with $|I\cap V_{L}|=|I\cap V_{R}|=k$. In general, each of the paths of length $12\ell+$ 1 in $G^{*}$ joining a vertex $u$ in $V_{L}$ to a vertex $v$ in $V_{R}$ is able to support an independent set of cardinality $\ell$ on its internal vertices (the vertices that are neither in $V_{L}$ nor in $V_{R}$). However, if $u$ and $v$ are both in $I$ then the internal vertices can only support an independent set of size $\ell-1$.
@@
The maximum of this expression is achieved at $k=n/(2+\alpha)$ where it takes the value $(3\ell+2/(2+\alpha))n$. When $k>\frac{2}{3}n$ there is a simpler argument. There are $3k$ edges of $G$ leaving the set $I\cap V_{L}$ and at most $3(n-k)$ of these can be accommodated in $V_{R}\setminus I$. So there are at least $6k-3n$ deficient paths and the cardinality of an independent set in this range is at most $2k+3\ell n-(6k-3n)=(3\ell+3)n-4k$. This expression is maximised at $k=\frac{2}{3}n$, at which point it evaluates to $(3\ell+\frac{1}{3})n$. Over the whole range of $k$, the maximum cardinality of an independent set in $\mathcal{I}_{=}$ is bounded by $(3\ell+2/(2+\alpha))n$.
@@
Corollary 18.
Excluding any finite set of connected graphs that includes at least one subdivided claw or path yields optimal mixing. Excluding any finite set of connected graphs none of which is a subdivided claw or path retains exponential time mixing.
Proof.
The positive direction is clear. For the negative direction, let $\mathcal{H}$ be the set of excluded graphs. For each $H\in\mathcal{H}$ there is a fugacity $\lambda_{H}$ and a degree of stretching $\ell_{H}$ such that Theorem 17 yields an infinite family of graphs on which Glauber dynamics requires exponential time for mixing. Now take $\lambda=\max_{H}\lambda_{H}$ and $\ell=\max_{H}\ell_{H}$. ∎

### v2

@@ 5. $S_{2,2,2}$-free graphs and beyond
First assume that $T$ has a single vertex $w$ such that either $w\not=v$ and $\deg(w)=4$ or $w=v$ and $\deg(w)=3$. Note that all vertices of $T$ other than $v,w,a,b,c$ have degree 2. By adding $e$ to $E(T)$ and removing the edges of $T$ lying between $w$ and $u_{j+1}$ we would obtain a smaller tree, contradicting minimality of $T$. (Notice that we must remove at least two edges.) So such an edge $e$ does not exist. By choice of $a,b,c$ we know that $d_{T}(w,a),d_{T}(w,b),d_{T}(w,c)\geq t+1$, so $G$ contains an induced $S_{t,t,t}$. (Note that we are measuring graph distances within the tree $T$ here.)
Next assume that $T$ has two vertices $w$ and $w^{\prime}$, with $w^{\prime}$ in a higher numbered level than $w$, and such that either (i) $w\not=v$ and and $\deg(w)=\deg(w^{\prime})=3$ or (ii) $w=v$ and $\deg(w)=2$ and $\deg(w^{\prime})=3$. Note that all vertices of $T$ other than $v,w,w^{\prime},a,b,c$ have degree 2. Suppose, without loss of generality, that $a$ and $b$ are descendants of $w^{\prime}$. With two exceptions, we may add $e$ to $E(T)$ and remove all the edges between $u_{j+1}$ and $w$, or $w^{\prime}$ as appropriate, to obtain a smaller tree, contradicting minimality of $T$. Those exceptions occur when $w^{\prime}\in L_{j}$, i.e., $u_{j}$ is in the same level as $w^{\prime}$. In that case, one edge is added and one removed and the size of the tree remains unchanged. There are two candidate vertices for $u_{j+1}$; these are the vertices in $L_{j+1}\cap V(T)$, say $x$ and $y$, that are adjacent to $w^{\prime}$. Let $x$ (respectively, $y$) be the vertex on the path in $T$ from $w^{\prime}$ to $a$ (respectively $b$). To summarise: The graph induced by $V(T)$ is tree $T$, with the possible addition of one or two of the edges $\{u_{j},x\}$ and $\{u_{j},y\}$. (Refer to Figure 5.)
@@
By choice of $a,b,c$ we know that $d_{T}(w^{\prime},a),d_{T}(w^{\prime},b)\geq t+1$. In the case that neither of the exceptional edges is present, we have an induced $S_{t,t,t}$ with its centre at $w^{\prime}$ and with the path from $w^{\prime}$ to $c$ routing via $w$. In the case of one exceptional edge, say $\{u_{j},x\}$, we have an induced $S_{t,t,t}$ with its centre at $x$ and with the path from $x$ to $b$ routing via $w^{\prime}$, and the path from $x$ to $c$ routing via $u_{j}$. In the case of two exceptional edges, we have an induced $S_{t,t,t}$ with its centre at $u_{j}$ and with the path from $u_{j}$ to $a$ routing via $x$, and the path from $u_{j}$ to $b$ routing via $y$.
@@
For any $\Delta\geq 3$, $\lambda>0$ and $t\geq 2$, Glauber dynamics for the hard-core distribution at fugacity $\lambda$ has optimal (i.e., $O(n\log n)$) mixing time on the class of $S_{t,t,t}$-free graphs of maximum degree $\Delta$.
@@ 6. Exponential-time mixing of $H$-free graphs
Let $H$ be a connected graph that is neither a subdivided claw nor a path. There exists a fugacity $\lambda$ and an infinite family $\mathcal{G}$ of $H$-free graphs of maximum degree 3 such that Glauber dynamics for the hard-core distribution on these graphs has exponential mixing time on $\mathcal{G}$.
@@
Let $\alpha>0$ be sufficiently small. Given $n$, let $G$ be a cubic bipartite graph on $n+n$ vertices that is a $1+\alpha$ expander for sets up to size $\frac{2}{3}n$. By this we mean the following. Let $V_{L}\cup V_{R}=V(G)$ be the bipartition of the vertex set of $G$. Then, every subset $S\subset V_{L}$ of size at most $\frac{2}{3}n$ is adjacent to at least $(1+\alpha)|S|$ vertices in $V_{R}$, and the same is true with $V_{L}$ and $V_{R}$ interchanged. Such graphs exist for some fixed $\alpha>0$ by a lemma due to Bassalygo [7]; see Alon [3, Lemma 4.1]. Let $\ell$ be a number and let $G^{*}$ denote the $(2\ell+1)$-stretch of $G$; that is, $G^{*}$ is obtained from $G$ by subdividing each edge by $2\ell$ new vertices. Fix $\ell$ sufficiently large that $G^{*}$ contains no induced $H$. Let $V_{L}$ and $V_{R}$ continue to denote the vertices in $G^{*}$ that are inherited from $G$.
@@
We will show that the independent sets in $\mathcal{I}_{=}$ all have cardinality significantly less than $(3\ell+1)n$. So if $\lambda$ is sufficiently large, the total weight $\mathrm{wt}(\mathcal{I}_{=})$ of independent sets in $\mathcal{I}_{=}$ will be exponentially smaller than $\mathrm{wt}(\mathcal{I}_{<})$. So consider an independent set $I$ with $|I\cap V_{L}|=|I\cap V_{R}|=k$. In general, each of the paths of length $12\ell+$ 1 in $G^{*}$ joining a vertex $u$ in $V_{L}$ to a vertex $v$ in $V_{R}$ is able to support an independent set of cardinality $\ell$ on its internal vertices (the vertices that are neither in $V_{L}$ nor in $V_{R}$). However, if $u$ and $v$ are both in $I$ then the internal vertices can only support an independent set of size $\ell-1$.
@@
The maximum of this expression is achieved at $k=n/(2+\alpha)$ where it takes the value $(3\ell+2/(2+\alpha))n$. When $k>\frac{2}{3}n$ there is a simpler argument. There are $3k$ edges of $G$ leaving the set $I\cap V_{L}$ and at most $3(n-k)$ of these can be accommodated in $V_{R}\setminus I$. So there are at least $6k-3n$ deficient paths and the cardinality of an independent set In $G^{*}$ in this range is at most $2k+3\ell n-(6k-3n)=(3\ell+3)n-4k$. This expression is maximised at $k=\frac{2}{3}n$, at which point it evaluates to $(3\ell+\frac{1}{3})n$. Over the whole range of $k$, the maximum cardinality of an independent set in $\mathcal{I}_{=}$ is bounded by $(3\ell+2/(2+\alpha))n$.
@@
Proof of Theorem 1.
The positive direction is immediate from Theorem 17. Now choose $\ell$ sufficiently large such that the stretched graph $G^{*}$ in the proof of Theorem 18 contains none of the graphs $H\in\mathcal{H}$. The construction of Theorem 18 yields a fugacity $\lambda$ and an infinite family of graphs (both depending on $\ell$) on which Glauber dynamics requires exponential mixing time at fugacity $\lambda$. ∎